OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y < -x + 4? (1, 2) (2, 1) (4, 0) (3, -2)

OpenStudy (anonymous):

(4, 0)

OpenStudy (anonymous):

0<-4+4 this is not true.

OpenStudy (anonymous):

Can you do another?

OpenStudy (anonymous):

yes,welcome.

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 6 > 13 or x + 6< 11

OpenStudy (anonymous):

where is the rest of question?

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 6 > 13 or x + 6 11 x > 7 or x 5 x 5 or x > 7 A number line with an open circle on 7, a closed circle on 5, and shading in between. A number line with an open circle on 7, shading to the right, and a closed circle on 5, shading to the left.

OpenStudy (anonymous):

what's that: x 5

OpenStudy (anonymous):

1st one Greater than or equal to 2nd is less than or equal to

OpenStudy (anonymous):

x + 6 > 13 x > 7 x + 6 less than or greater to 11 x is less than or greater to 5 the answer is: A number line with an open circle on 7, a closed circle on 5, and shading in between.

OpenStudy (anonymous):

Thnak you Life saver

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Last One i Promise :) Which of the following is not an equivalent form of the compound inequality x + 12 > 20 and x + 12 26 20 < x + 12 26 A number line with a closed circle on 14, shading to the right, and an open circle on 8, shading to the left. 8 < x and x 14 A number line with a closed circle on 14, an open circle on 8, and shading in between.

OpenStudy (anonymous):

x + 12 > 20 and x + 12 <= 26 x > 20-12 and x <= 26-12 x > 8 and x <= 14 so the answer is A number line with a closed circle on 14, shading to the right, and an open circle on 8, shading to the left.